A market’s fixed basket of Goods: 100 Tennis Balls; 200 Golf Balls
YEAR PRICE OF TENNIS BALL PRICE OF GOLF BALL
2017 (base year) $2 $4
2018 $3 $6
2019 $4 $7

a) Compute the cost of the basket of goods in each year, the consumer price index in each year and the inflation rate for the year 2018 and 2019.
b) Write the three differences between CPI and GDP deflator.
1
Expert's answer
2020-04-13T09:42:54-0400

a) cost of basket: 2017 = $1,000.

2018 = $1,500.

2019 = $1,800.

"CPI2018 = 1500\/1000*100=150%." %.

"CPI2019 = 1800\/1000*100=180" %.

Inflation:

2018 = "(150-100)\/1000*100=50" %.

2019 = "(180-150)150*100=20" %.

b)The first difference is that the GDP deflator measures the prices of all goods and services produced, whereas the CPI or RPI measures the prices of only the goods and services bought by consumers. 

The second difference is that the GDP deflator includes only those goods produced domestically. 

The third difference concerns how the two measures aggregate the many prices in the economy. The CPI is computed using a fixed basket of goods, whereas the GDP deflator allows the basket of goods to change over time as the composition of GDP changes.